5. If P(E) = 0.05, what is the probability of ‘not E’?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: We know that, P(E)+P(not E) = 1 It is given that, P(E) = 0.05 So, P(not E) = 1-P(E) Or, P(not E) = 1-0.05 ∴ P(not E) = 0.95